a electric iron draws a current of 0.5A when voltage is 200V . calculate the charge flowing through it in one hour.
1
Expert's answer
2012-05-31T08:07:55-0400
Charge flowing through the iron& is calculated using the following formula Q = I·T
Substituting given values we obtain
Q = 0.5*3600 = 1800 C.